OpenStudy (anonymous):

Sarah’s mother has a miniature village modeled after her home town. The model is 10 feet long and 7 feet wide. The actual village is 7 miles long. What is the village’s actual width? If necessary, round to the nearest tenth of a mile 4.9 miles 10 miles 1 mile 0.7 mile which one is the answer?

OpenStudy (anonymous):

yay u r on

OpenStudy (anonymous):

This is a ratio problem. The model is 10 feet long and the village is 7 miles long. If you compare those, you get the ratio of the model size to the real thing: 10 feet / 7 miles That is the same ratio as the model's width the real village width: 7 feet / ??? miles

OpenStudy (anonymous):

so: 10 feet / 7 miles = 7 feet / x miles Then you can cross-multiply to simplify: 10x = 49 And then divide both sides by 10 to solve for x. x = 49/10

OpenStudy (tkhunny):

\(\dfrac{7 mile}{10 ft} = \dfrac{7 mile\cdot\dfrac{5280 ft}{1 mile}}{10 ft} = \dfrac{36960}{10}\) Can you finish?

OpenStudy (anonymous):

so wouldnty it be 4.9 miles because the ratio is 49/10

OpenStudy (anonymous):

sorry, that drawing was all messed up :(

OpenStudy (anonymous):

its ok tht drawing thing is weird

OpenStudy (anonymous):

yes, the ratio of lengths is 10 feet to 7 miles, and the equivalent ratio of widths is 7 feet to 4.9 miles. To actually get the ratio, you would need to convert miles to feet, but all you really need to do is to find the equivalent ratio and solve for the unknown width.

OpenStudy (anonymous):

thx lifesaver

OpenStudy (anonymous):

You're welcome, happy to help. All these problems with models or shrinking or growing something are similar ratio problems...

OpenStudy (anonymous):

oh hey before i go is this right x/2 = 8/9 x = 0.44

OpenStudy (anonymous):

huh? :) Is this a different problem?

OpenStudy (anonymous):

yes yes

OpenStudy (anonymous):

So, what is the question?

OpenStudy (anonymous):

its a proportion x/2 = 8/9

OpenStudy (anonymous):

would x = 0.44

OpenStudy (anonymous):

Oh, ok :) Yes, you had an extra "x" in it earlier... x/2 = 8/9 Multiply both sides by 2: x = 16/9 Then you can use a calculator or whatever if you need the decimal. x = 1.7778 I think you divided both sides by 2 instead of multiplying.

OpenStudy (anonymous):

x/2 = 8/9 The other way is to cross multiply: 9x = (2)(8) Simplify: 9x = 16 Divide both sides by 9 to solve for x: x = 16/9

OpenStudy (anonymous):

thx so much

OpenStudy (anonymous):

ur so awesome

OpenStudy (anonymous):

:)